OpenStudy (anonymous):

Find all the real cube roots of 0.000027

OpenStudy (anonymous):

0.000027=(27/1000000)=(3/100)^3 Hence the real cube root is 3/100 .@pineapple2

OpenStudy (anonymous):

@pineapple2